OpenStudy (anonymous):

In how many ways can four hearts be drawn from a deck of cards? (Order is not important.) I know it is a combination and that the answer is 715, but I just need help on how to get that answer.

geerky42 (geerky42):

Ok, so you have to choose 4 hearts from total of 13 hearts in deck of cards, right?\[\binom {13} {4}=\dfrac{13!}{4!(13-4)!} =\dfrac{13!}{4!~9!} =\boxed{715}\]

OpenStudy (anonymous):

Thank you so much! It makes sense now!

geerky42 (geerky42):

No problem.